An empty bulb has a volume of 350. mL and a mass of 82.561 g. When filled with an unknown gas at 733 mmHg and 22 0C, the bulb has a mass of 82.786 g. Is the gas argon (Ar), methane (CH4), or nitrogen monoxide (NO)?
1
Expert's answer
2017-12-18T15:41:07-0500
pV=m/MRT, M=mRT/Pv
T=273+22=295K, V=0.000350m³
R=8/314J/(Mol*K)
m=82.786-82.561=0.225g
p=97725.35Pa
M(CH₄)=16g/Mol